OpenStudy (lukecrayonz):

Which are natural numbers, integers, rational numbers, irrational numbers and real numbers? 0, 1, 1/12, 1/16, 1/20

OpenStudy (lukecrayonz):

natural, 1 integers, 0,1

OpenStudy (lukecrayonz):

and all are real

OpenStudy (unklerhaukus):

all are rational

OpenStudy (unklerhaukus):

which is a rational number

OpenStudy (lukecrayonz):

Wait, yes, sorry, was thinking of something else

OpenStudy (anonymous):

\[ 0, 1, \frac 1{12}, \frac 1{16}, \frac 1{20} \] There is no imaginary part. Hence all are reals. They are also rationals since each of them can be written in the form of \(\frac pq \) where \(p\) and \(q\) are integers and \(q\neq 0\)

OpenStudy (anonymous):

The set of Naturals usually start from 1, but according to some convention we also consider 0 to be natural. Please check which one is expedient for you.

OpenStudy (anonymous):

Here only 0 and 1 are integers.

OpenStudy (anonymous):

None, is irrational since the set of rational and irrationals set are disjoint. REF: An irrational number is any real number that cannot be expressed as a ratio a/b, where a and b are integers, with b non-zero, and is therefore not a rational number. (Source: http://en.wikipedia.org/wiki/Irrational_number )

OpenStudy (unklerhaukus):

b canot be one.